12kg of carbon burns fully in 64g of pure oxygen amd the volume of products is measured at 273k and 1 atm the volume of gas is


1
Expert's answer
2021-02-16T07:45:10-0500

Given,

Mass of carbon ="12 kg"


Moles of Carbon="\dfrac{12}{12}=1000\text{mol}"


Mass of Oxygen =64g


Moles of oxygen ="\dfrac{64}{32}=2\text{mol}"


The Balance Chemical reaction is given by-

"C+O_2\rightarrow CO_2"


Here The limiting reagent is "O_2" ,


Thus, Mole of "CO_2" = Mole of "O_2=2\text{mol}"


Also, Pressure "P=1 \text{atm}"

Temprature, "T=273K"

Universal gas constant, "R=8.314Jk^-mol^-"


Let Volume be V

According to ideal gas equation,

"PV=nRT"

"1\times V=2\times 8.314\times273"

"V=4539.44ml=4.54l"


Hence Volume of product formed is "4.54l"